OpenStudy (anonymous):

Solve by the substitution method 2x-10=-y 3x-y=0

OpenStudy (anonymous):

Minus the 2x on the first equation to get -2x-y=-10 3x-y=0

OpenStudy (anonymous):

Lets get rid of the y so multiply a - on the 1st equation to get 2x+y=10 3x-y=0 Then it will be 5x=10 Divide the 5 on both sides x=2 plug the 2 on one of the equations 3(2)-y=0 6-y=0 Subtract the 6 -y=-6 Divide the - y=6 So it is x=2 & y =6

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Good use of the elimination method, but to answer the problem as posed, change the second equation to 3x=y and put it into the first equation 2x-10=-3x then 5x=10 then x=2 and y =6. Substitution method.

OpenStudy (anonymous):

(x,y)=(2,6) (almost forgot to give the answer!!)